OpenStudy (anonymous):

Evaluate the expression below for x = 1, y = –2, and z = 3. 3x-2y/z+4

OpenStudy (uri):

\[\left(\begin{matrix}3(1)-2(-2) \\ 3+4\end{matrix}\right)\]

OpenStudy (uri):

Can you do it now?

OpenStudy (anonymous):

so its -7/1

OpenStudy (uri):

-7/7 = -1

OpenStudy (anonymous):

thanks